Q1: For a solution in which µ = 5.0 * 10-2, calculate K'sp for
(a) AgSCN. (Ksp 1.1*10-12)
(b) La(IO3)3. (Ksp 1.1*10-11)
Q2: Use activities to calculate the molar solubility of Zn(OH)2 in (Ksp 3*10-16)
(a) 0.0100 M KCl.
(b) 0.0330 M K2SO4.
1
Expert's answer
2020-12-29T02:42:46-0500

Q1.

a) Ksp 1.1*10-12.

b) Ksp 1.1*10-11.

Q2.

a) 3*10-16 x 0.01 = 3*10-18.

b) 3*10-16 x 0.033 = 9,9*10-18